Operating Systems and Systems Integration

Is this a boring subject only about the theory of the

internal operation of an operating system???

OSSI — ver. 1.1 Introducton to OSSI — slide 1

Is this a theoretical subject?

• Hmmm, operating systems: sounds like a theoretical subject • At university, a lot of theory from a text book

• Are we going to spend lots of time copying from a text book? …

OSSI — ver. 1.1 Introducton to OSSI — slide 2

NO! This is practical!

• This subject aims to provide you with practical skills that you will find useful in your workplace.

• You learn here by doing.

• I even understand this subject too!

OSSI — ver. 1.1 Introducton to OSSI — slide 3

Mostly based on Linux

• … with some comparisons with Windows 2000.

• Aim to support career certification, e.g., Red Hat Certified Engi-

neer, Linux Professional Institute

OSSI — ver. 1.1 Introducton to OSSI — slide 4 OSSI — ver. 1.1 Introducton to OSSI — slide 5

So why does Nick love Linux?

• It is free software

• Free as in freedom

• … also, free as in free beer

• It works really well

• It can “glue” many other things together

• Like Lego; can build anything I want

OSSI — ver. 1.1 Introducton to OSSI — slide 6

More reasons why Nick Loves Linux

• Built on cooperation

• Great for Internet applications

• Only fully compliant tcp^{/}ip

• We can see and modify the source code to any of it that we want to

• Based on standards

• Uses open protocols

OSSI — ver. 1.1 Introducton to OSSI — slide 7

Open Standards and Protocols

• What do I mean by “Open Standard?”

∘ Can freely download the standard, not pay a huge fee just to

read it ∘ Agreed to by open discussion: barrier to participation is lack

of ability, not lack of money! ∘ Examples of open standards:

– tcp^{/}ip, http, ssh, ftp, dns, dhcp, Perl, ldap, smtp,

tls, many, many more

OSSI — ver. 1.1 Introducton to OSSI — slide 8

Standards that are not so open

• Many standards are developed in a less open way

• Example: the wep protocol, used to provide “privacy” for wireless lans

• Developed behind closed doors, announced to the public

• Cracked almost immediately: a bad design

• Not open to inspection and peer review

OSSI — ver. 1.1 Introducton to OSSI — slide 9

Embrace and Extend

• Many companies use “standards”

• But change them a little bit to give a competitive advantage • Example: Kerberos

∘ Developed at MIT as open source software, ∘ Improved by programmers all round the world ∘ Used by Microsoft for authentication in Active Directory

• Microsoft changed the interpretation of one small part of the protocol

• Effect: all Kerberos clients can use a Microsoft Kerberos server • But Microsoft clients will refuse to work with anything except a

Microsoft server.

OSSI — ver. 1.1 Introducton to OSSI — slide 10

Other “Industry Standards”

• Consider the Microsoft Office suite

• The layout is a secret

• OpenOffice.org has developed an office suite that can read and write Office documents

• Very hard work:

∘ a moving target

∘ Much effort by ms to make them very hard to read, and even

harder to write

OSSI — ver. 1.1 Introducton to OSSI — slide 11

What is “Systems Integration”?

• It involves combining products from many companies into a system

• Other words: interoperation, compatibility

• Very important: it’s not enough just to learn one product

OSSI — ver. 1.1 Introducton to OSSI — slide 12

Why From Many Companies?

• Why not just from one supplier?

OSSI — ver. 1.1 Introducton to OSSI — slide 13

Monoculture (One supplier)

• Advantages:

• Fewer system integration problems

• Less skill required

• All training from the one vendor

• Disadvantages:

• Vulnerability in one is a vulnerability in all (e.g., So.Big, Blaster, SirCam, NIMBDA worms, Outlook viruses)

• One supplier cannot make everything:

• E.g., Cisco sell more routers and switches than Microsoft

OSSI — ver. 1.1 Introducton to OSSI — slide 14

How Best to Integrate? Standards

• Using Open Standards that are free of patents and other restrictions on use

• … But every vendor says they support the standards!

• Some standards are more open than others!

OSSI — ver. 1.1 Introducton to OSSI — slide 15

So what answers are there?

• Free software supports open standards

• So open, you can read the source code and see how it works!

• You can even change it to suit your needs, and if you find problems, you can fix them!

• Aim for interoperation, not for exclusion or market benefit

OSSI — ver. 1.1 Introducton to OSSI — slide 16

Samba — Systems Integration

• Samba allows a Linux or Unix or Macintosh machine to talk with Windows

• and the other way round

• Free Software

• A Linux machine can be an NT-4 compatible Primary Domain Controller

• Can also be an AD member

• Very stable, high quality

• Used by many companies to interoperate with Windows

OSSI — ver. 1.1 Introducton to OSSI — slide 17

Practical: 60 hours of laboratory workshops!

• We start this week!

• So what will we do?

• In the first class, we partition hard disks

• … then install Linux.

• See you there!

OSSI — ver. 1.1 Introducton to OSSI — slide 18